# TDD Without Tears: Towards Test Case Generation from Requirements through Deep Reinforcement Learning

### v1

@@ 6. Related Works
In this section, we highlight the difference in our work with respect to the literature on reinforcement learning in software engineering.
@@
Reinforcement learning (RL) (Sutton and Barto, 2018) is a machine-learning approach where an agent learns to map observations to actions by maximizing numerical reward functions. Prior studies have demonstrated successes in various tasks ranging from recreation games (Silver et al., 2018; Berner et al., 2019; Jaderberg et al., 2019; Vinyals et al., 2019) to real-world nuclear fusion plasma control (Degrave et al., 2022). Recently, RL has been applied to text generation tasks. Specifically, LLMs utilize RL to train models based on user preferences through human feedback (RLHF) (Christiano et al., 2017). Consequently, LLMs can exhibit complex and coherent behaviors such as instruction following (Ouyang et al., 2022), summarization (Stiennon et al., 2020), stylistic continuation (Ziegler et al., 2019), and conversation (Jaques et al., 2019).
@@
In software engineering, reinforcement learning has been adopted in the code generation tasks (Wang et al., 2022; Le et al., 2022) to improve the compilability and execution rate of the generated code. For example, Wang et al. (Wang et al., 2022) proposed CompCoder, an RL approach that considers compiler feedback in the reward function to generate more compilable code. Le et al. (Le et al., 2022) proposed CodeRL, an actor-critic (Bahdanau et al., 2016; Konda and Tsitsiklis, 1999) deep reinforcement learning framework for code generation.
@@
Different from prior studies, we formulate the Text-to-Testcase task as a deep reinforcement learning problem. Results also show that our PyTester outperforms other LLMs, highlighting the novelty and the significance of our work that contributes to the area of Text-to-Testcase generation.
7. Threat to Validity
In this section, we disclose the threat to the validity of our study.
@@ 7. Threat to Validity
Threats to internal validity. The variation of the reward function may impact the performance of PyTester. However, our RQ2 confirms that, among the 7 variations, the performance varies between 57.31% and 69.87% of the TestCaseScore. Despite the high difference in the performance among the reward functions, we find that the lowest-performing variation of our PyTester still outperforms LLMs. Thus, the reward function does not pose a threat to the validity of our study. Nevertheless, our study is limited by one scheme of reward points (ranging from -2 to 4). However, finding an optimal scheme of reward points is beyond the scope of this work. Thus, future research is encouraged to explore the impact of the reward points on the performance of Deep RL models.
Threats to external validity. The experiment of this paper is limited to the APPS benchmark dataset, the Python programming language, and the context of the online programming tasks. Thus, the results may not be generalized to other datasets, programming languages, and contexts. Thus, future research is encouraged to explore the generalizability of our PyTester in other contexts.
8. Conclusion
@@
  • Of the test cases generated by PyTester, 99% are syntactically correct, 84% are aligned with the requirements while achieving a code coverage of 80% and a mutation score of 61%, which outperforms all of the studied LLMs.
@@
  • The SE domain knowledge should be considered when designing a deep reinforcement learning architecture. The performance of many AI models often relies on its model architecture. Thus, many prior studies often adopted various data mining and optimization techniques to improve such AI models. Different from existing work, this paper takes an alternative direction by integrating the SE domain knowledge when designing the reinforcement learning architecture (i.e., incorporating test case characteristics into the reward function). This enables PyTester to generate test cases that are more syntactically correct, executable, complete, and effective while being aligned with a given requirement.

### v2

@@ 7 Related Works
In this section, we highlight the difference in our work to the literature on reinforcement learning in software engineering.
@@
Reinforcement learning (RL) (Sutton and Barto, 2018) is a machine-learning approach where an agent learns to map observations to actions by maximizing numerical reward functions. Recently, RL has been applied to text generation tasks. Specifically, LLMs utilize RL to train models based on user preferences through human feedback (RLHF) (Christiano et al., 2017). Consequently, LLMs can exhibit complex and coherent behaviors such as instruction following (Ouyang et al., 2022), summarization (Stiennon et al., 2020), and conversation (Jaques et al., 2019).
@@
In software engineering, reinforcement learning has been adopted in the code generation tasks (Wang et al., 2022; Le et al., 2022) to improve the compilability and execution rate of the generated code. For example, Wang et al. (2022) proposed CompCoder, an RL approach that considers compiler feedback in the reward function to generate more compilable code. Le et al. (2022) proposed CodeRL, an actor-critic (Bahdanau et al., 2016; Konda and Tsitsiklis, 1999) deep reinforcement learning framework for code generation.
@@
Different from prior studies, we formulate the Text-to-Testcase task as a deep reinforcement learning problem. We incorporate the domain knowledge such as test case characteristics (e.g., syntax and code coverage scores) into the model through a reward function. Additionally, we perform a deep analysis of ablation studies on reward components, showing the importance of the test case characteristics. Finally, the results also show that our PyTester outperforms other LLMs, highlighting the novelty and the significance of our work that contributes to the area of Text-to-Testcase generation.
8 Threat to Validity
In this section, we disclose our threat to the validity.
@@ 8 Threat to Validity
Data leakage is often a critical concern when applying large language models (LLMs) in software engineering (i.e., the phenomenon when data samples in the testing dataset were included in the training dataset). With PyTester, it is based on the CodeT5-large model, which is pre-trained on the CodeSearchNet dataset, and then later fine-tuned on the APPS benchmark dataset. Prior to the pre-training and fine-tuning step, we carefully analyzed and ensured that none of the testing samples in the APPS dataset appeared in the CodeSearchNet (Husain et al., 2019) and the training set of the APPS dataset to avoid any data leakage problem. Thus, we confirm that the results of our PyTester will not be impacted by the data leakage. With GPT-3.5 (baseline comparison), we are unable to determine if there is any data leakage during the model training of GPT-3.5 due to the close-source nature of GPT-3.5. Despite the risk of data leakage with the GPT-3.5 model, our experiments still confirm that PyTester (with no risk of data leakage) outperforms GPT-3.5 (with the risk of data leakage) for all evaluation measures.
Threats to internal validity. The variation of the reward function may impact the performance of PyTester. However, our RQ2 confirms that, among the 7 variations, the performance varies between 57.31% and 69.87% of the TestCaseScore. Despite the high difference in the performance among the reward functions, we find that the lowest-performing variation of our PyTester still outperforms most of the LLMs. Thus, the reward function does not pose a threat to the validity of our study. Nevertheless, our study is limited by one scheme of reward points (ranging from -2 to 4). However, finding an optimal scheme of reward points is beyond the scope of this work. Thus, future research is encouraged to explore the impact of the reward points on the performance of Deep RL models.
The primary goal of software testing is to identify defects and vulnerabilities within the software. Consequently, automated test case generation is expected to produce test cases that can find bugs accurately. To evaluate this bug-finding capability of a model, a dataset containing buggy code is typically required. However, due to the unavailability of the bug information in our APPS benchmark dataset, it is challenging to directly assess the bug-finding capability of PyTester. To address this challenge, we include the evaluation of the mutation testing to our paper. Mutation testing works by introducing small bugs into the correct code (e.g., replacing a minus operator with a plus operator). If the test cases still pass the mutated code, they will receive a zero Mutation Score, indicating that the test cases fail to identify bugs in the mutated code. This metric provides valuable insight into how well and how sensitive the generated test cases can identify bugs, serving as an alternative means of assessing the potential capability of a model in bug-finding.
Threats to External Validity. We acknowledge that our PyTester is currently limited to the generated test cases based on the textual description of one single call. In real-world scenarios, test cases are more complex (including test prefixes, etc). Thus, future research may consider the generation of more complex test cases that include test prefixes, etc.
PyTester training dataset is limited to the APPS benchmark dataset—the only Text-to-Testcase benchmark dataset with sufficient size. In addition to broadening our analysis and ensuring the performance of PyTester, we provide additional evaluations in the zero-shot setting on the HumanEval and MBPP datasets. Future research may consider increasing the generalizability of our PyTester approach with other datasets, programming languages, and contexts.
9 Conclusion
@@
  • Of the test cases generated by PyTester, 99% are syntactically correct, 84% are passed while achieving a code coverage of 80% and a mutation score of 61%, which outperforms all of the studied LLMs.
@@
  • The SE domain knowledge should be considered when designing a deep reinforcement learning architecture. The performance of many AI models often relies on its model architecture. Thus, many prior studies often adopted various data mining and optimization techniques to improve such AI models. Different from existing work, this paper takes an alternative direction by integrating the SE domain knowledge when designing the reinforcement learning architecture (i.e., incorporating test case characteristics into the reward function). This enables PyTester to generate test cases that are more syntactically correct, executable, complete, and effective while being aligned with a given description.